# NEURAL EMBEDDING COMPRESSION FOR EFFICIENT MULTI-TASK EARTH OBSERVATION MODELLING

### v1

@@ 1 INTRODUCTION
To overcome these issues, we propose the exchange of EO-FM embeddings computed at the data provider instead of the raw data. It follows that novel data only has to be processed once by the computationally expensive encoder The resulting embeddings can be stored and reused for multiple downstream applications. The inference can be performed at the data consumer with a lightweight head.
@@
Our contributions in this paper are threefold: i) The discussion on EO-FM embedding sharing compared to raw data, ii) the conceptual integration of neural compression into the data pipeline, and iii) the benchmarking of the neural compression approach for various downstream applications.
@@ 2 RELATED WORK
Learned neural compression was pioneered by Balle et al. [11] with their seminal work on integrating information theoretical concepts into Deep Neural Network training for data compression. A variety of works followed proposing more powerful entropy models [12, 15] and extending its applications to video compression [14, 16, 13] and even neural network compression [17]. Recently, some works have leveraged masked auto-encoding (MAE) as a lossy compression method, dropping data at the image patch [18] or embedding level [19].
CORSA [20] trains a VQVAE [21] for the purposes of compression for satellite imagery. In their work, they briefly explore training an MLP directly on the VQVAE embeddings for a simple binary classification task, but their method does not show conclusive results. To the best of our knowledge, ours is the first work that demonstrates the effectiveness of the direct use of compressed embeddings in downstream tasks by combining FMs with neural compression. We target EO applications as input data is often sourced through queries on large repositories, whereas other application areas may rely on user input. This characteristic makes it a good fit for NEC.
@@ 3 METHOD
In this section, we frame our method as the optimization of a rate-distortion objective following the framework of end-to-end learned compression established in Balle et al. [11]. Traditionally, these methods aim to transmit an input image or video using the smallest amount of bits possible (R - Rate) while minimizing the loss in quality (D - Distortion) after reconstruction by a decoder network. This trade-off is controlled by $\lambda$, resulting in an optimization problem of the form $\min\lambda R+D$.
@@
Specifically, we use MAE [23] with a standard vision transformer (ViT) [24] architecture. MAE masks part of the input and tasks the model with recovering it based on the unmasked input. This is achieved by minimizing the mean squared error between the masked input and its reconstruction, which we take as our distortion term. We choose MAE due to its simplicity and computational efficiency and the ViT-B architecture due to its wide use in recent EO models [5, 6, 3]. However, any loss that encourages generally useful embeddings for EO and any compatible encoder architecture can be used.
Rate Term: Balle et al. [11] establishes the general methodology for computing the rate term in the full loss. We give a short overview and direct the reader to [12, Appendix 6.1] for more details.
Given an embedding vector $y$, the minimum number of bits required to transmit $y$ is given by Shannon’s source coding theorem [25] as $-\log_{2}p(y)$, where $p$ defines the probability distribution over all embedding vectors. Our goal is to minimize the expectation of this quantity over the distribution of all embedding vectors $-\mathbb{E}\left[\log_{2}p(y)\right]$, which is precisely the entropy of this distribution. We adapt the fully factorized entropy model defined in Balle et al. [12] in order to model the probability distribution of these embeddings for MAE. Specifically, we leverage the implementation in the CompressAI [26] library.
@@ 5 Results
1. Scene Classification: Fig. 3(a) plots the results for the UCM Dataset. For the high-performance case, we observe that UQE performs well until $5$ bit quantization and is only slightly outperformed by NEC. For higher compression levels, this does not hold. NEC is able to achieve $94\%$ accuracy at an average embedding size of $0.47$KB, a reduction of only $5\%$ in accuracy for a size reduction of $\sim$700700700700x, greatly outperforming 3 and 2-bit UQE. The benefit of transmitting the first layer of the decoder is also apparent, particularly as the compression level increases.
@@ 6 CONCLUSION
In this work, we introduce NEC, a framework for downstream training and inference for EO tasks based on directly transmitting compressed embeddings. We show how any FM can be adapted to produce compressed embeddings of different fidelities using a single architecture and with minimal compute. Our evaluation concludes with a data transfer reduction of one to three orders of magnitude with only limited degradation in performance.

### v3

@@ 1 INTRODUCTION
To overcome these issues, we propose the exchange of EO-FM embeddings computed at the data provider instead of the raw data. It follows that novel data only has to be processed once by the computationally expensive encoder. The resulting embeddings can be stored and reused for multiple downstream applications. The inference can be performed at the data consumer with a lightweight head.
@@
Our contributions in this paper are threefold: i) The discussion on EO-FM embedding sharing compared to raw data, ii) the integration of neural compression into the data pipeline for classification and segmentation tasks, and iii) the benchmarking of the neural compression approach on two downstream EO applications.
@@ 2 RELATED WORK
Learned neural compression, popularized by Ballé et al. [11], aims to integrate information theoretical concepts into neural network training for data compression. A variety of works followed proposing more powerful entropy models [12, 15] and extending its applications to video compression [14, 16, 13] and even neural network compression [17].
Singh et al. [18] explore learned compression for embeddings, but produce task-specific encoders. Dubois et al. [19] lay out the theoretical foundation for learned compression of predictive features and adapt CLIP [20] to produce compressed embeddings for generic classification tasks. We expand this scope to explore compressed embeddings for pixel-wise prediction as well as classification. We target EO applications as input data is often sourced from queries on large remote repositories, rather than from local user input, making it a good fit for NEC.
@@ 3 METHOD
In this section, we frame our method as the optimization of a rate-distortion objective following the framework of end-to-end learned compression established in Ballé et al. [11]. Traditionally, these methods aim to transmit an input image or video using the smallest amount of bits possible (R - Rate) while minimizing the loss in quality (D - Distortion) after reconstruction by a decoder network. This trade-off is controlled by $\lambda$, resulting in an optimization problem of the form $\min\lambda R+D$.
@@
Specifically, we use MAE [22] with a standard vision transformer (ViT) [23] architecture. MAE masks part of the input and tasks the model with recovering it based on the unmasked input. This is achieved by minimizing the mean squared error between the masked input and its reconstruction, which we take as our distortion term. We choose MAE due to its simplicity and applicability to a large range of downstream tasks and the ViT-B architecture due to its wide use in recent EO models [5, 6, 3]. However, any loss that encourages generally useful embeddings for EO and any compatible encoder architecture can be used.
Rate Term: Ballé et al. [11] establishes the general methodology for computing the rate term in the full loss. We give a short overview and direct the reader to [12, Appendix 6.1] for more details.
Given an embedding vector $y$, the minimum number of bits required to transmit $y$ is given by Shannon’s source coding theorem [24] as $-\log_{2}p(y)$, where $p$ defines the probability distribution over all embedding vectors. Our goal is to minimize the expectation of this quantity over the distribution of all embedding vectors $-\mathbb{E}\left[\log_{2}p(y)\right]$, which is precisely the entropy of this distribution. We adapt the fully factorized entropy model defined in Ballé et al. [12] in order to model the probability distribution of these embeddings for MAE. Specifically, we leverage the implementation in the CompressAI [25] library.
@@ 5 Results
1. Scene Classification: Fig. 2(a) plots the results for the UCM Dataset. For the high-performance case, we observe that UQE performs well until $5$ bit quantization and is only slightly outperformed by NEC. For higher compression levels, this does not hold. NEC is able to achieve $94\%$ accuracy at an average embedding size of $0.47$KB, a reduction of only $5\%$ in accuracy for a size reduction of $\sim$700700700700x, greatly outperforming 3 and 2-bit UQE. In particular, transmitting the first layer of the decoder is crucial to retain performance at very high compression.
@@ 6 CONCLUSION
In this work, we introduce NEC, a framework for downstream training and inference for EO tasks based on directly transmitting compressed embeddings. We show how any FM can be adapted to produce compressed embeddings of different fidelities using a single architecture and with minimal compute. Our evaluation demonstrates a data transfer reduction of one to three orders of magnitude with only limited degradation in performance.